OpenStudy (anonymous):

Osmotic pressure of a 4.44% solution of anhydrous CaCl2 was found to be 16.42 atm at 27°C. What is the degree of dissociation of CaCl2 ?

OpenStudy (mayankdevnani):

4.44% (w/V) of CaCl2

OpenStudy (mayankdevnani):

\[\large \bf \pi=cRT\] where, \[\bf \pi=osmotic~pressure,R=Universal~gas~constant,T=temp.(K)\]

OpenStudy (mayankdevnani):

\[\large \bf c=\frac{no.of~moles~of CaCl_2}{Vol.of~Solution}\]

OpenStudy (mayankdevnani):

but ,you have to determine degree of dissociation,so \[\large \bf \pi=i \times cRT\] where, \[\large \bf i=van't~hoff~factor\]

OpenStudy (mayankdevnani):

solve for `i`

OpenStudy (mayankdevnani):

and we know that \[\large \bf i=1+(n-1)\alpha~~~~~~~\rightarrow For~Dissociation\]

OpenStudy (mayankdevnani):

where, n=3

OpenStudy (mayankdevnani):

and solve for `alpha` you will get your answer !

OpenStudy (mayankdevnani):

hope you understand. @Tushi